OpenStudy (babyslapmafro):

At t=0sec a ball, initially at rest, starts to roll down a ramp with constant acceleration. You notice it moves 1ft between t=0sec and t=1sec. How far does it move between t=1 and t=2sec?

OpenStudy (anonymous):

xo=vo*t-(.5)at^2 (a common kinematic equation)

OpenStudy (anonymous):

xo=displacement vo=initial velocity (in your case zero) t=time a=acceleration which you want to find.

OpenStudy (babyslapmafro):

Ok, so my acceleration is -2ft/sec^2

OpenStudy (babyslapmafro):

@cherio12

OpenStudy (anonymous):

positive, xo would be negative in this case because it is going down

OpenStudy (anonymous):

but yet, 2m/s^2

OpenStudy (anonymous):

yes*

OpenStudy (babyslapmafro):

ok so now I plug the acceleration back into the initial equation to find displacement?

OpenStudy (anonymous):

yes, .5*2*(2^2)=x-1

OpenStudy (babyslapmafro):

I'm just not entirely clear why the acceleration is positive and the displacement is negative

OpenStudy (anonymous):

there are the same sign

OpenStudy (anonymous):

in the equation i wrote -.5*a*t^2

OpenStudy (babyslapmafro):

if they are the same sign the result be a negative final displacement

OpenStudy (anonymous):

it is the common formula, but it depends on the direction the acceleration is acting. If it acts the same way the object goes its positive

OpenStudy (anonymous):

like...you could argue that the acceleration of the ball is DOWN the hill. and the ball roles DOWN the hill

OpenStudy (babyslapmafro):

Hmm...I used a positive displacement and negative acceleration

OpenStudy (babyslapmafro):

kinda confused

OpenStudy (anonymous):

well, so why do you believe the acceleration is negative?

OpenStudy (babyslapmafro):

because it is downward acceleration

OpenStudy (anonymous):

okay i would agree. so now why do you believe displacement is positive?

OpenStudy (babyslapmafro):

because it is traveling a positive distance away from a starting point

OpenStudy (babyslapmafro):

on a horizontal axis

OpenStudy (anonymous):